# LaM-SLidE: Latent Space Modeling of Spatial Dynamical Systems via Linked Entities

### v3

@@ 2 $\mathbf{o}\mathrel{+}=\text{ParallelMLPAttentionWithRoPE}(\mathbf{o},\bm{\tau},\text{dim}=1)$
Using einops [76] notation, the latent layer in Figure 6 can be expressed as:
$\mathbf{o}^{\prime}$ $\leftarrow\texttt{rearrange}(\mathbf{o},\;\texttt{($B$ $L$) $T$ $D_{z}$}\rightarrow\texttt{($B$ $T$) $L$ $D_{z}$})$
@@
$\mathbf{o}^{\prime}$ $\leftarrow\texttt{rearrange}(\mathbf{o}^{\prime},\;\texttt{($B$ $T$) $L$ $D_{z}$}\rightarrow\texttt{($B$ $L$) $T$ $D_{z}$})$
@@
with parameters sets $\psi$ and $\phi$, where for the latent block the time dimension gets absorbed into the batch dimension and for the temporal block the latent dimension gets absorbed into the batch dimension.
@@ Appendix C Additional Information on Stochastic Interpolants
The stochastic interpolants framework  [2, 3, 60] is defined without reference to an forward SDE, which allows a lot of flexibility, any choice of $\alpha_{t}$ an $\sigma_{t}$, satisfying the following conditions is possible:
@@ D.1 Molecular Dynamics (MD)
[…]
Analytic solutions of $\psi$ for specific operators $\hat{H(t)}$ are hardly known and are only available for simple systems like free particles or hydrogen atoms. In contrast to that are proteins with many thousands of atoms. However, already for much smaller quantum systems approximations are needed. A famous example is the Born–Oppenheimer approximation, where the wave function of the multi-body system is decomposed into parts for heavier atom nuclei and the light-weight electrons, which usually move much faster. In this case, one obtains a Schrödinger equation for electron movement and another Schrödinger equation for nuclei movement. A much faster option than solving a second Schrödinger equation for the motion of the nuclei is to use the laws from classical Newtonian dynamics. The solution of the first Schrödinger equation defines an energy potential, which can be utilized to obtain forces $\mathbf{F}_{i}$ on the nuclei and to update nuclei positions according to Newton’s equation of motion: $\mathbf{F}_{i}=m_{i}\ \ddot{\mathbf{q}}_{i}(t)$ (with $m_{i}$ being the mass of particle $i$ and $\mathbf{q}_{i}(t)$ describing the motion trajectory of particle $i$ over time $t$).
Additional complexity in studying molecule dynamics is introduced by environmental conditions surrounding molecules. Maybe the most important is temperature. For bio-molecules it is often of interest to assume that they are dissolved in water. To model temperature, a usual strategy is to assume a system of coupled harmonic oscillators to model a heat bath, from which Langevin dynamics can be derived [29, 101]. The investigation of the relationship between quantum-mechanical modeling of heat baths and Langevin dynamics still seems to be a current research topic, where there there are different aspects like the coupling of the oscillators or Markovian properties when stochastic forces are introduced. For instance, Hoel & Szepessy [39], studies how canonical quantum observables are approximated by molecular dynamics. This includes the definition of density operators, which behave according to the quantum Liouville-von Neumann equation.
@@
Newton’s equations of motions for all particles under consideration form a system of ordinary differential equations (ODEs), to which different numeric integration schemes like Euler, Leapfrog, or, Verlet can be applied to obtain particle position trajectories for given initial positions and initial velocities. In case temperature is included, the resulting Langevin equations form a system of stochastic differential equations (SDEs), and Langevin integrators can be used. It should be mentioned, that it is often necessary to use very small integration timesteps to avoid large approximation errors. This, however, increases the time needed to find new stable molecular configurations.
@@ D.2 Relationship to Graph Foundation Models
From our perspective, LaM-SLidE bears a relationship to graph foundation models [GFMs; 56, 62]. Bommasani et al. [16] consider foundation models to be trained on broad data at scale and to be adaptable to a wide range of downstream tasks. Mao et al. [62] argue, that graphs are more diverse than natural language or images, and therefore there are quite unique challenges for GFMs. Especially they mention that none of the current GFM have the capability to transfer across all graph tasks and datasets from all domains. It is for sure true that LaM-SLidE is not a GFM in this sense. However, it might be debatable whether LaM-SLidE might serve as a domain- or task-specfic GFM. While we mainly focused on a trajectory prediction task and are from that point of view task-specific, we observed that our trained models can generalize across different molecules or differently taken scenes, which might seem quite remarkable given that it is common practice to train specific trajectory prediction models for single molecules or single scenes. Nevertheless, it was not our aim in this research to provide a GFM, since we believe that this would require more investigation into further domains and could also require, for instance, checking whether emergent abilities might arise with larger models and more training data [56].
@@ D.3 Relationship to Video and Language Diffusion Models
In video diffusion models, the model unmasks future frames; in language diffusion models, the model unmasks unknown tokens. Both paradigms learn to recover information that is initially obscured in the sequence, and importantly, both methods do that in parallel over the whole input sequence [9], compared to autoregressive models which predict a single frame or a single token at a time.
Similarly, LaM-SLidE represents each timestep as a set of latent tokens (or alternatively, as a single token when concatenated). This perspective allows us to seamlessly incorporate recent advances from both video and language diffusion research into our modeling paradigm.
@@ E.1 Datasets
The pedestrian movement dataset with data processing is available at https://github.com/MediaBrain-SJTU/EqMotion.
@@
The dataset creation scripts, along with its predefined splits is available at https://github.com/hanjq17/GeoTDM.
@@
The MD17 dataset is available at http://www.sgdml.org/#datasets. Preprocessing and dataset splits follow Han et al. [35] and can be accessed through their GitHub repository at https://github.com/hanjq17/GeoTDM. The dataset comprises, 5,000 training, 1000 validation and 1000 test trajectories for each molecule.
@@
The dataset, including the full simulation parameters for ground truth simulations, is sourced from Jing et al. [44] and is publicly available in their GitHub repository at https://github.com/bjing2016/mdgen. The dataset comprises 3,109 training, 100 validation and 100 test peptides.
@@ E.2 Condition and Prediction Horizon
Table 7 shows the conditioning and prediction horizon for the individual experiments. For the Tetrapeptides experiments we predicted 1000 steps in parallel and reconditioned the model ten times on the last frame for each predicted block, this concept is similar to Arriola et al. [9].
@@ E.3 Implementation Details
(i) First Stage. In the first stage we train the encoding and decoding functions $\mathcal{E}$ and $\mathcal{D}$ in an auto-encoding fashion, i.e., we optimize for a precise reconstruction of the original system state representation from its latent representation well. For discrete features (e.g., atom type, residue type) we tend to use a cross-entropy loss, whereas for continuous features we use a regression loss (e.g., position, distance). The loss functions for each individual task are summarized in Section E.5. Notably, also the entity identifier assignment is random. (ii) Second Stage. In the second stage, we freeze the encoder and train the approximator to model the temporal dynamics via the encoded latent system representations. To learn a consistent behavior over time, we pass $\mathbf{U}$ from the encoder $\mathcal{E}$ to the decoder $\mathcal{D}$. To avoid high variance latent spaces we used layer-normalization [10] (see. Section E.3).
@@
For the embedding of the identifiers we use a torch.nn.Embedding [66] layer, where we assign a random subset of the possible embeddings to the entities in each training step. See also Algorithm 1.
@@
To avoid high variance latent spaces, Rombach et al. [77] relies on KL-reg., imposing a small KL-penalty towards a standard normal on the latent space, as used in VAE [49]. Recent work [99] has shown that layer normalization [10] can achieve similar regulatory effects without requiring an additional loss term and simplifying training procedure, we adapt this approach in our method (see left part of Figure 2).
@@
For the latent Flow Model we additionally apply auxiliary losses for the individual tasks, as shown in Section E.5. Where we decode the the predicted latent system representations and back-propagate through the frozen decoder to the latent model.
@@
For the experiments on tetrapeptides in Section 4.5, we employ the Atom14 representation as used in AlphaFold [1]. In this representation, each entity corresponds to one amino acid of the tetrapeptide, where multiple atomic positions are encoded into a single vector of dimension $D_{x}=3\times 14$. Masked atomic positions are excluded from gradient computation during model updates. This representation is computationally more efficient.
@@ E.4 Loss Functions
$\mathcal{L}_{\mathrm{pos}}(\mathbf{X}^{t},\hat{\bf{X}}^{t})=\frac{1}{N}\sum_{i=1}^{N}||\mathbf{X}_{i}^{t}-\hat{\mathbf{X}}_{i}^{t}||_{2}^{2}$ (13)
@@
Depending on the experiment we have different CE losses depending on the problem see Section E.5.
@@ E.5 Hyperparameters
Tabs. 8, 9, 10, 11 and 12 show the hyperparameters for the individual tasks, loss functions are as defined in Section E.4. For all trained models we use the AdamW [50, 58] optimizer and use EMA [31] in each update step with a decay parameter of $\beta=0.999$.
@@ E.6 Evaluation Details
Our analysis of the Tetrapeptide trajectories utilized PyEMMA [84] and followed the procedure as Jing et al. [44], incorporating both Time-lagged Independent Component Analysis (TICA) [70] and Markov State Models (MSM) [42]. For the evaluation of these metrics we relied on the implementations provided by [44].
@@ E.8 Software
Table 8: Hyperparameter configuration for the pedestrian movement experiments (Section 4.1).
First Stage Network Encoder Number of latents $L$ 2 Number of entity embeddings 8 Number of attention heads 2 Number of cross attention layers 1 Dimension latents $D_{z}$ 32 Dimension entity embedding 128 Dimension attention head 16 Decoder Number of attention heads 2 Number of cross attention layers 1 Dimension attention head 16 Loss Weight $\mathcal{L}_{pos}(\mathbf{X},\hat{\mathbf{X}})$ 1 $\mathcal{L}_{int}(\mathbf{X},\hat{\mathbf{X}})$ 1 Training Learning rate 1e-4 Learning rate scheduler CosineAnnealing(min_lr=1e-7) Batch size 256 Epochs 3K Precision 32-Full Batch size 1024 Second Stage Setup Condition 8 Frames Prediction 12 Frames Network Hidden dimension 128 Number of Layers 6 Auxiliary - Loss Weight $\mathcal{L}_{pos}(\mathbf{X},\hat{\mathbf{X}})$ 0.25 $\mathcal{L}_{int}(\mathbf{X},\hat{\mathbf{X}})$ 0.25 Training Learning rate 1e-3 Batch size 64 Epochs 1K Precision BF16-Mixed Inference Integrator Euler ODE steps 10
Table 9: Hyperparameter configuration for the basketball player movement experiments (Section 4.2).
First Stage Network Encoder Number of latents $L$ 32 Number of entity embeddings 11 Number of attention heads 2 Number of cross attention layers 1 Dimension latents $D_{z}$ 32 Dimension entity embedding 128 Dimension attention head 16 Decoder Latent dimension 32 Number of attention heads 8 Number of cross attention layers 1 Loss Weight $\mathcal{L}_{pos}(\mathbf{X},\hat{\mathbf{X}})$ 1 $\mathcal{L}_{int}(\mathbf{X},\hat{\mathbf{X}})$ 1 $\mathcal{L}_{CE}(\cdot,\cdot)\,-\text{Group }$ 0.01 $\mathcal{L}_{CE}(\cdot,\cdot)\,-\text{Team}$ 0.01 Training Learning rate 1e-4 Learning rate scheduler CosineAnnealing(min_lr=1e-7) Optimizer AdamW Batch size 16 Second Stage Setup Condition 8 Frames Prediction 12 Frames Network Hidden dimension $H$ 128 Number of Layers 6 Auxiliary - Loss Weight $\mathcal{L}_{pos}(\mathbf{X},\hat{\mathbf{X}})$ 0.25 $\mathcal{L}_{int}(\mathbf{X},\hat{\mathbf{X}})$ 0.25 Training Learning rate 1e-3 Learning rate scheduler CosineAnnealing(min_lr=1e-7) Batch size 64 Epochs 500 Precision BF16-Mixed Inference Integrator Euler ODE steps 10
Table 10: Hyperparameter configuration for the N-Body experiments (Section 4.3).
First Stage Network Encoder Number of latents $L$ 16 Number of entity embeddings 10 Number of attention heads 2 Number of cross attention layers 1 Dimension latents $D_{z}$ 32 Dimension entity embedding 128 Dimension attention head 16 Decoder Number of cross attention layers 1 Number of attention heads 2 Number of cross attention layers 16 Loss Weight $\mathcal{L}_{pos}(\mathbf{X},\hat{\mathbf{X}})$ 1 $\mathcal{L}_{int}(\mathbf{X},\hat{\mathbf{X}})$ 1 Training Learning rate 1e-3 Batch size 128 Epochs 2K Precision 32-Full Second Stage Setup Condition 10 Frames Prediction 20 Frames Network Hidden dimension 256 Number of Layers 6 Auxiliary - Loss Weight $\mathcal{L}_{pos}(\mathbf{X},\hat{\mathbf{X}})$ 0.0 $\mathcal{L}_{int}(\mathbf{X},\hat{\mathbf{X}})$ 0.0 Training Learning rate 1e-3 Learning rate scheduler CosineAnnealing(min_lr=1e-7) Batch size 64 Epochs 1K Precision BF16-Mixed Inference Integrator Euler ODE steps 10
Table 11: Hyperparameter configuration for the small molecule (MD17) experiments (Section 4.4).
First Stage Network Encoder Number of latents $L$ 32 Number of entity embeddings 8 Number of attention heads 2 Number of cross attention layers 1 Dimension latents $D_{z}$ 32 Dimension entity embedding 128 Dimension attention head 16 Decoder Number of cross attention layers 1 Number of attention heads 2 Number of cross attention layers 16 Loss Weight $\mathcal{L}_{pos}(\mathbf{X},\hat{\mathbf{X}})$ 1 $\mathcal{L}_{int}(\mathbf{X},\hat{\mathbf{X}})$ 1 $\mathcal{L}_{CE}(\cdot,\cdot)\,-\text{ Atom type}$ 1 Training Learning rate 1e-4 Batch size 256 Epochs 3K Precision 32-Full Second Stage Setup Condition 10 Frames Prediction 20 Frames Network Hidden dimension 128 Number of Layers 6 Auxiliary - Loss Weight $\mathcal{L}_{pos}(\mathbf{X},\hat{\mathbf{X}})$ 0.25 $\mathcal{L}_{int}(\mathbf{X},\hat{\mathbf{X}})$ 0.25 Training Learning rate 1e-3 Learning rate scheduler CosineAnnealing(min_lr=1e-7) Batch size 64 Epochs 2K Precision BF16-Mixed Inference Integrator Euler ODE steps 10
Table 12: Hyperparameter configuration for the Tetrapeptides experiments (Section 4.5).
First Stage Network Encoder Number of latents $L$ 5 Number of entity embeddings 8 Number of attention heads 2 Number of cross attention layers 1 Dimension latents $D_{z}$ 96 Dimension entity embedding 128 Dimension attention head 16 Decoder Number of attention heads 2 Number of cross attention layers 1 Dimension attention head 16 Loss Weight $\mathcal{L}_{pos}(\mathbf{X},\hat{\mathbf{X}})$ 1 $\mathcal{L}_{int}(\mathbf{X},\hat{\mathbf{X}})$ 1 $\mathcal{L}_{frame}(\mathbf{X},\hat{\mathbf{X}})$ 1 $\mathcal{L}_{tors}(\mathbf{X},\hat{\mathbf{X}})$ 0.1 $\mathcal{L}_{CE}(\cdot,\cdot)\,-\text{Residue type}$ 0.001 Training Learning rate 1e-4 Batch size 16 Epochs 200K Precision 32-Full Second Stage Setup Condition 1 Frame Prediction 10,000 Frames (10x rollouts) Network Hidden dimension 384 Number of Layers 6 Auxiliary - Loss Weight $\mathcal{L}_{pos}(\mathbf{X},\hat{\mathbf{X}})$ 0.25 $\mathcal{L}_{int}(\mathbf{X},\hat{\mathbf{X}})$ 0.25 $\mathcal{L}_{frame}(\mathbf{X},\hat{\mathbf{X}})$ 0.25 Training Learning rate 1e-3 Optimizer AdamW Batch size 64 Epochs 1.5K Precision BF16-Mixed Inference Integrator Dopri5 [20] ODE steps adaptive

### v4

@@ 2 $\mathbf{o}\mathrel{+}=\text{ParallelMLPAttentionWithRoPE}(\mathbf{o},\bm{\tau},\text{dim}=1)$
Using einops [76] notation, the latent layer in Figure˜6 can be expressed as:
[…]
@@
[…]
@@
with parameters sets $\psi$ and $\phi$, where for the latent block, the time dimension gets absorbed into the batch dimension, and for the temporal block, the latent dimension gets absorbed into the batch dimension.
B.5 Python Pseudocode
This section shows Python pseudocode for training and inference for the latent approximator model.
## Latent layer of the approximator (high level)
z = z + embed_cond(z_cond)
t_embed = embed_time(t)
# T x L x Dz, attention is calculated over L | ref [4,5]
z = attention_spatial(z, t_embed)
z = rearrange(z, "T L Dz -> L T Dz")
# L x T x Dz, attention is calculated over T | ref [4,5]
z = attention_temporal(z, t_embed)
z = rearrange(z, "L T Dz -> T L Dz")
return z
### Training ####
# T x N x Dn (N=number of entities, D=[x,y,z,atom_type])
# e.g. Aspirin 21x4
x
ids
# T binary mask [1, 0, 0, ...], e.g., one conditioning frame
mask
# Encoding and conditioning setup
# T x L x Dz Encoder state into latent vector
z1 = encode(x, ids)
# T x L x Dz Contains only the conditioning frames,
# the other frames are masked out and set to 0.
z_cond = mask_frames(z1, mask)
# Sample z0 from a normal distribtion
z0 = torch.randn_like(z1)
# Sample t uniformly
t = torch.uniform(0, 1)
# Create interpolation
zt = t * z1 + (1 - t) * z0
# Loss
z1_hat = approximator(zt, t, z_cond)
loss = torch.mean((z1_hat - z1)**2)
loss.backward()
### Inference ####
# T x N x Dn, but non-conditioning frames are zero
x_cond
# Transform model trained on data prediction into velocity
# prediction model.
velocity_model = data_to_velocity_model(approximator)
# T x L x Dz, contains only the conditioning
# frames/ missing frames are set to 0
z_cond = encode(x_cond, ids)
z0 = torch.randn_like(z1)
# Solve ODE torchdiffeq (https://github.com/rtqichen/torchdiffeq)
z1_pred = solver(velocity_model, z0, z_cond)
x_hat = decoder(z1_pred)
@@ Appendix C Additional Information on Stochastic Interpolants
The stochastic interpolants framework  [2, 3, 60] is defined without reference to a forward SDE, which allows a lot of flexibility; any choice of $\alpha_{t}$ and $\sigma_{t}$, satisfying the following conditions, is possible:
@@ D.1 Molecular Dynamics (MD)
The most fundamental concepts used to describe the dynamics of molecules nowadays are those given by the laws of quantum mechanics. The Schrödinger equation is a partial differential equation that gives the evolution of the complex-valued wave function $\psi$ over time $t$: $i\hbar\dfrac{\partial\psi}{\partial t}=\hat{H}(t)\psi$. Here $i$ is the imaginary unit with $i^{2}=-1$, $\hbar$ is the reduced Planck constant, and $\hat{H}(t)$ is the Hamiltonian operator at time $t$, which is applied to a function $\psi$ and maps to another function. It determines how a quantum system evolves with time, and its eigenvalues correspond to measurable energy values of the quantum system. The solution to Schrödinger’s equation in the many-body case (particles $1,\ldots,N$) is the wave function $\psi(\mathbf{x}_{1},\ldots,\mathbf{x}_{N},t):\bigtimes_{i=1}^{N}\mathbb{R}^{3}\times\mathbb{R}\rightarrow\mathbb{C}$ which we abbreviate as $\psi(\left\{\mathbf{x}\right\},t)$. It’s the square modulus $|\psi(\left\{\mathbf{x}\right\},t)|^{2}=\psi^{*}(\left\{\mathbf{x}\right\},t)\psi(\left\{\mathbf{x}\right\},t)$ is usually interpreted as a probability density to measure the positions $\mathbf{x}_{1},\ldots,\mathbf{x}_{N}$ at time $t$, whereby the normalization condition $\int\ldots\int|\psi(\left\{\mathbf{x}\right\},t)|^{2}d\mathbf{x}_{1}\ldots d\mathbf{x}_{N}=1$ holds for the wave function $\psi$.
Analytic solutions of $\psi$ for specific operators $\hat{H(t)}$ are hardly known and are only available for simple systems like free particles or hydrogen atoms. In contrast to that are proteins, which consist of many thousands of atoms. However, already for much smaller quantum systems, approximations are needed. A famous example is the Born–Oppenheimer approximation, where the wave function of the multi-body system is decomposed into parts for heavier atom nuclei and the light-weight electrons, which usually move much faster. In this case, one obtains a Schrödinger equation for the electron’s movement and another Schrödinger equation for the nucleus’s movement. A much faster option than solving a second Schrödinger equation for the motion of the nuclei is to use the laws from classical Newtonian dynamics. The solution of the first Schrödinger equation defines an energy potential, which can be utilized to obtain forces $\mathbf{F}_{i}$ on the nuclei and to update nuclei positions according to Newton’s equation of motion: $\mathbf{F}_{i}=m_{i}\ \ddot{\mathbf{q}}_{i}(t)$ (with $m_{i}$ being the mass of particle $i$ and $\mathbf{q}_{i}(t)$ describing the motion trajectory of particle $i$ over time $t$).
Additional complexity in studying molecule dynamics is introduced by the environmental conditions surrounding molecules. Maybe the most important property is temperature. For biomolecules, it is often of interest to assume that they are dissolved in water. To model temperature, a usual strategy is to assume a system of coupled harmonic oscillators to model a heat bath, from which Langevin dynamics can be derived [29, 102]. The investigation of the relationship between quantum-mechanical modeling of heat baths and Langevin dynamics remains a current research topic, with various aspects, including the coupling of oscillators and the introduction of Markovian properties with stochastic forces. For instance, Hoel & Szepessy [39] studies how canonical quantum observables are approximated by molecular dynamics. This includes the definition of density operators, which behave according to the quantum Liouville-von Neumann equation.
@@
Newton’s equations of motion for all particles under consideration form a system of ordinary differential equations (ODEs), to which various numerical integration schemes, such as Euler, Leapfrog, or Verlet, can be applied to obtain particle position trajectories for given initial positions and velocities. If temperature is included, the resulting Langevin equations form a system of stochastic differential equations (SDEs), and Langevin integrators can be employed. It should be mentioned that it is often necessary to use very small integration timesteps to avoid large approximation errors. This, however, increases the time needed to find new stable molecular configurations.
@@ D.2 Relationship to Graph Foundation Models
From our perspective, LaM-SLidE bears a relationship to graph foundation models [GFMs; 56, 62]. Bommasani et al. [16] consider foundation models to be trained on broad data at scale and to be adaptable to a wide range of downstream tasks. Mao et al. [62] argue that graphs are more diverse than natural language or images, and therefore, there are unique challenges for GFMs. Especially, they mention that none of the current GFM have the capability to transfer across all graph tasks and datasets from all domains. It is for sure true that LaM-SLidE is not a GFM in this sense. However, it might be debatable whether LaM-SLidE might serve as a domain- or task-specific GFM. While we primarily focused on a trajectory prediction task and are, from that point of view, task-specific, we observed that our trained models can generalize across different molecules or scenes, which may seem quite remarkable given that it is common practice to train specific trajectory prediction models for individual molecules or scenes. Nevertheless, it was not our aim in this research to provide a GFM, as we believe that this would require further investigation into additional domains and could also necessitate, for instance, examining whether emergent abilities might arise with larger models and more training data [56].
@@ D.3 Relationship to Video and Language Diffusion Models
In video diffusion models, the model unmasks future frames; in language diffusion models, the model unmasks unknown tokens. Both paradigms learn to recover information that is initially obscured in the sequence, and importantly, both methods do so in parallel over the entire input sequence [9], compared to autoregressive models, which predict a single frame or token at a time.
Similarly, LaM-SLidE represents each timestep as a set of latent tokens (or as a single token when concatenated). This perspective allows us to seamlessly incorporate recent advances from both video and language diffusion research into our modeling paradigm.
@@ E.1 Datasets
The pedestrian movement dataset, along with its data processing, is available at https://github.com/MediaBrain-SJTU/EqMotion.
@@
The dataset creation scripts, along with their predefined splits, are available at https://github.com/hanjq17/GeoTDM.
@@
The MD17 dataset is available at http://www.sgdml.org/#datasets. Preprocessing and dataset splits follow Han et al. [35] and can be accessed through their GitHub repository at https://github.com/hanjq17/GeoTDM. The dataset comprises 5,000 training, 1000 validation, and 1000 test trajectories for each molecule.
@@
The dataset, including the full simulation parameters for ground truth simulations, is sourced from Jing et al. [44] and is publicly available in their GitHub repository at https://github.com/bjing2016/mdgen. The dataset comprises 3,109 training, 100 validation, and 100 test peptides.
@@ E.2 Condition and Prediction Horizon
Table˜7 shows the conditioning and prediction horizon for the individual experiments. For the Tetrapeptides experiments, we predicted 1000 steps in parallel and reconditioned the model ten times on the last frame for each predicted block, this concept is similar to Arriola et al. [9].
@@ E.3 Implementation Details
(i) First Stage. In the first stage, we train the encoding and decoding functions $\mathcal{E}$ and $\mathcal{D}$ in an auto-encoding fashion, i.e., we optimize for a precise reconstruction of the original system state representation from its latent representation. For discrete features (e.g., atom type, residue type), we tend to use a cross-entropy loss, whereas for continuous features we use a regression loss (e.g., position, distance). The loss functions for each task are summarized in Section˜E.5. Notably, the entity identifier assignment is also random. (ii) Second Stage. In the second stage, we freeze the encoder and train the approximator to model the temporal dynamics via the encoded latent system representations. To learn a consistent behavior over time, we pass $\mathbf{U}$ from the encoder $\mathcal{E}$ to the decoder $\mathcal{D}$. To avoid high variance latent spaces, we used layer-normalization [10] (see Section˜E.3).
@@
For the embedding of the identifiers we use a torch.nn.Embedding [66] layer, where we assign a random subset of the possible embeddings to the entities in each training step. See also Algorithm˜1.
@@
To avoid high variance latent spaces, Rombach et al. [77] relies on KL-reg., imposing a small KL-penalty towards a standard normal on the latent space, as used in VAE [49]. Recent work [100] has shown that layer normalization [10] can achieve similar regulatory effects without requiring an additional loss term and simplifying the training procedure. We adapt this approach in our method (see the left part of Figure˜2).
@@
For the latent Flow Model we additionally apply auxiliary losses for the individual tasks, as shown in Section˜E.5. Where we decode the predicted latent system representations and back-propagate through the frozen decoder to the latent model.
@@
For the experiments on tetrapeptides in Section˜4.5, we employ the Atom14 representation as used in AlphaFold [1]. In this representation, each entity corresponds to one amino acid of the tetrapeptide, where multiple atomic positions are encoded into a single vector of dimension $D_{x}=3\times 14$. Masked atomic positions are excluded from gradient computation during model updates. This representation is computationally more efficient.
@@ E.4 Loss Functions
[…]
@@
Depending on the experiment, we have different CE losses depending on the problem, see Section˜E.5.
@@ E.5 Hyperparameters
Tabs.˜8, 9, 10, 11 and 12 show the hyperparameters for the individual tasks, loss functions are as defined in Section˜E.4. For all trained models, we use the AdamW [50, 58] optimizer and use EMA [31] in each update step with a decay parameter of $\beta=0.999$.
@@ E.6 Evaluation Details
Our analysis of the Tetrapeptide trajectories utilized PyEMMA [84] and followed the procedure as Jing et al. [44], incorporating both Time-lagged Independent Component Analysis (TICA) [70] and Markov State Models (MSM) [42]. For the evaluation of these metrics, we relied on the implementations provided by [44].
@@ E.8 Software
Table 8: Hyperparameter configuration for the pedestrian movement experiments (Section˜4.1).
[…]
Table 9: Hyperparameter configuration for the basketball player movement experiments (Section˜4.2).
[…]
Table 10: Hyperparameter configuration for the N-Body experiments (Section˜4.3).
[…]
Table 11: Hyperparameter configuration for the small molecule (MD17) experiments (Section˜4.4).
[…]
Table 12: Hyperparameter configuration for the Tetrapeptides experiments (Section˜4.5).
[…]